# CGD: Modifying the Loss Landscape by Gradient Regularization

### v2

@@ 1 Introduction
Line-search methods such as gradient descent and Newton’s method have been extremely popular in solving unconstrained optimization problems [15]. However, vanilla version of such algorithms typically produce unsatisfactory results when applied to large scale optimization problems such as those involving deep neural networks. In such problems, the objective/loss function has a very non-linear landscape that is embedded with several local maxima and minima making them difficult to optimize [11]. To tackle this problem, one of the approaches used is to modify the objective/loss function in such a way that the new function is easier to optimize. This is typically done by adding a regularization term that will make the loss landscape more smooth, and which would in turn reduce the chances of the optimizer reaching sub-optimal minima, thereby improving generalizability [5, 18, 1, 17, 8].
@@
Towards this, we propose Constraint gradient descent (CGD), a variant of the GD algorithm which achieves this by constraining the gradient norm to be appropriately small. Instead of solving this constrained optimization problem, we consider the Lagrangian of this function and perform GD on it. The Lagrangian parameter $\lambda$ controls the penalty on gradient norm and thereby controls the steepness of the modified function. It is on this modified function that we seek to apply GD to possibly attain iterates that are much closer to the local minima as compared to GD iterates. From Fig. 1 (Right) we additionally see that compared to GD, the CGD iterates (which is nothing but GD iterates on $f_{2}$) have a lower gradient norm. Since the path to minima is over points with lower gradient norm, this we believe is an attractive feature and even amounts to better generalization properties for high dimensional functions such as loss landscapes of neural networks. The idea of flat minima and their attractiveness goes long back to Hochreiter and Schmidhuber [7]: An optimal minimum is considered “flat” if the test error changes less in its neighbourhood. Keskar et al. [10] and Chaudhari et al. [4] observe better generalization results for neural networks at flat minima. Gradient regularization has only been recently explored, that too from a numerical perspective for deep neural networks (DNNs) where it has been shown that fixed learning rates result in implicit gradient regularization  [17, 8, 1, 16].
@@
  • We propose a new line-search procedure called Constraint Gradient Descent (CGD) that performs gradient descent on a gradient regularized loss function. We also provide global linear convergence for CGD in Theorem 3.1.
@@ 2.1 Line-Search Methods
We focus on the situation where ${\mathbf{x}}^{*}$ is obtained using an iterative line-search procedure (for details refer [12, Ch. 3]). The steepest descent (gradient descent) method, Newton’s method, and quasi-Newton’s method are some examples of line-search methods. In an iterative algorithm, the key idea is to begin with an initial guess $\mathbf{x}_{0}$ for the minimizer and generate a sequence of vectors $\mathbf{x}_{0},\mathbf{x}_{1},\ldots$ until convergence (or until desired level of accuracy is achieved). In such algorithms, $\mathbf{x}_{k+1}$ is obtained using $\mathbf{x}_{k}$ using a pre-defined update rule.
@@ 2.2 Quasi-Newton Methods
We consider DFP and BFGS algorithms that update $G_{k}$ using symmetric rank-two updates at each descent step [3, 6]. The update equations for DFP and BFGS are given as follows (for more details refer [12, Ch. 6]):
[…]
@@
We can derive the corresponding Hessian approximation $\tilde{G}_{k}$ from the update steps for $G_{k}$ by applying the Sherman-Morrison-Woodbury formula [12, App. A]. These are given as follows:
[…]
@@ 3 Constrained Gradient descent
Let $\mathcal{S}_{\hat{\mathbf{x}}}$ and $\mathcal{S}_{\mathbf{x}^{\star}}$ be the stationary points of $f({\mathbf{x}})$ and $g({\mathbf{x}})$ as defined in Equation 8. Then, $\mathcal{S}_{\hat{{\mathbf{x}}}}\subseteq\mathcal{S}_{{\mathbf{x}}^{*}}$. Furthermore, for any ${\mathbf{x}}^{*}\in\mathcal{S}_{{\mathbf{x}}^{*}}$ one of the following is true: (a) ${\mathbf{x}}^{*}\in\mathcal{S}_{\hat{{\mathbf{x}}}}$ or (b) $\nabla f({\mathbf{x}}^{*})$ is an eigenvector of $H({\mathbf{x}}^{*})$ with the eigenvalue $-\frac{1}{2\lambda}$.
@@
As a result, descent on the penalized function might actually cause ascent over the original loss function. For example in Fig. 2, observe that at $x=-7.7$ (marked with the dashed line), steepest descent along $-\nabla g({\mathbf{x}})$ direction would actually cause an ascent over the original loss function.
To fix this behaviour, we ensure that we only move along the direction $-\nabla g({\mathbf{x}})$ when it is a descent direction. Therefore, we only move along $-\nabla g({\mathbf{x}})$, if $\nabla f({\mathbf{x}})^{T}\nabla g({\mathbf{x}})>0$ (from Equation 2). Otherwise, we set $\lambda=0$ at this point and move along the $-\nabla f({\mathbf{x}})$ direction. Note that this simple check also helps us to avoid stopping at artificially introduced stationary points. We summarize CGD in Algorithm 1.
@@
2:${\mathbf{x}}_{k+1}\leftarrow{\mathbf{x}}_{k}+\alpha{\mathbf{p}}_{k}$Else
@@ 3.1 Global Linear Convergence of CGD
[…]
@@ 3.2 CGD-FD: Finite difference approximation of the Hessian
3:${\mathbf{p}}_{k}\leftarrow-(1-\nu)\nabla f_{k}-\nu\nabla f({\mathbf{x}}_{k}+r\nabla f_{k})$
@@
5:${\mathbf{x}}_{k+1}\leftarrow{\mathbf{x}}_{k}+\alpha{\mathbf{p}}_{k}$Else
6:${\mathbf{x}}_{k+1}\leftarrow{\mathbf{x}}_{k}-\alpha\nabla f_{k}$
@@ 3.3 CGD-QN: Quasi-Newton Variants of CGD
We consider quasi-Newton variants of CGD (CGD-QN) where a positive-definite approximation of the Hessian is maintained. The update step for CGD-QN is given as: ${\mathbf{x}}_{k+1}={\mathbf{x}}_{k}-\alpha\left(I_{n}+2\lambda_{k}\tilde{G}_{k}\right)\nabla f_{k}$ where $\tilde{G}_{k}$ is the Hessian approximation at step $k$, which is updated after every step. Particularly Equations 4 and 5 correspond to the updates of DFP and BFGS variants of CGD respectively.
@@
3:${\mathbf{x}}_{k+1}\leftarrow{\mathbf{x}}_{k}+\alpha{\mathbf{p}}_{k}$Else
4:${\mathbf{x}}_{k+1}\leftarrow{\mathbf{x}}_{k}-\alpha\nabla f_{k}$EndIf
@@ 4 Numerical Experiments
We test CGD-FD and CGD-QN on synthetic test functions from Virtual Library of Simulation Experiments: Test Functions and Datasets  [^1] http://www.sfu.ca/~ssurjano.
@@
Quadratic function $10$ 0.4 0.01 18.89 97.91
Rotated hyper-ellipsoid function $5$ 0.5 0.01 15.94 82.76
Levy function $2$ L(0.01, 0.1) 0.05 23.73 63.21
Branin function $2$ 0.07 0.01 37.53 87.07
Griewank function $2$ 40.0 0.01 0.01 0.08
Matyas function $2$ 10.0 0.01 1.83 34.40
@@
Table 1 summarizes our findings and Fig. 3 visualizes function values vs gradient evaluations for the chosen test functions for experiments on CGD-FD. Observe that in the first step of optimizing Quadratic function (Fig. 3a), CGD-FD has a $97.91\%$ decrease in value compared to that of GD which only had an initial improvement of about $18.89\%$. Similar pattern can also be seen for other functions. This gives us the insight that CGD-FD penalizes the original function well enough to concentrate most of the decrease in value within the first few steps of the trajectory itself. Also, notice that in Griewank function (Fig. 3e) even though the initial step improvement is low ($=0.08\%$), we still see a much rapid decrease in value within the first 10 steps compared to that of the GD trajectory. Thus, using CGD-FD with appropriate penalization can provide for great boosts in function value well within the starting iterates of the optimization procedure.
Experiments for CGD-QN
: We ran CGD-BFGS and CGD-DFP methods against BFGS and DFP for $T=40$. The function values vs iterations are visualized in Fig. 4. We observe that for most suitable choices of $\lambda$, CGD-DFP and CGD-BFGS exhibit similar behaviour and hence, their trajectories coincide. Another thing to observe is that in CGD-QN methods, most improvements aren’t concentrated in the initial steps since Hessian approximation is still not very great here. Instead the improvements are more gradual and appear over further steps. For example, in Zakharov function (Fig. 4a), we can observe that all methods follow the same decrease in function value for the first 1–2 steps and then the CGD-QN methods start to drop more in value. In EggHolder function (Fig. 4c) this effect is most pronounced where the improvements are very gradual but providing with better results than their vanilla counterparts.

### v3

@@ 1 Introduction
3D surface plots showing the "original loss function f₁(x)" (a blue, shallow curve) and the "steeper loss function f₂(x)" (a brown, steeper curve).
A contour plot and gradient norm heatmap of the original loss function f₁(x) overlaid with two 5-step GD trajectories, one over f₁ and the other over f₂. Both start from the same point, but the path over f₂ reaches the local minimum more directly.
@@
Line-search methods such as gradient descent and Newton’s method have been extremely popular in solving unconstrained optimization problems [17]. However, vanilla version of such algorithms typically produce unsatisfactory results when applied to large scale optimization problems such as those involving deep neural networks. In such problems, the objective/loss function has a very non-linear landscape that is embedded with several local maxima and minima making them difficult to optimize [12]. To tackle this problem, one of the approaches used is to modify the objective/loss function in such a way that the new function is easier to optimize. This is typically done by adding a regularization term that will make the loss landscape more smooth, and which would in turn reduce the chances of the optimizer reaching sub-optimal minima, thereby improving generalizability [7, 5, 23, 1, 22, 9].
@@
Towards this, we propose Constrained gradient descent (CGD), a variant of the GD algorithm which achieves this by constraining the gradient norm to be appropriately small. Instead of solving this constrained optimization problem, we consider the Lagrangian of this function and perform GD on it. The Lagrangian parameter $\lambda$ controls the penalty on gradient norm and thereby controls the steepness of the modified function. It is on this modified function that we seek to apply GD to possibly attain iterates that are much closer to the local minima as compared to GD iterates. From Fig. 1 (Right) we additionally see that compared to GD, the CGD iterates (which is nothing but GD iterates on $f_{2}$) have a lower gradient norm. Since the path to minima is over points with lower gradient norm, this we believe is an attractive feature and even amounts to better generalization properties for high dimensional functions such as loss landscapes of neural networks. The idea of flat minima and their attractiveness goes long back to Hochreiter and Schmidhuber [8]: An optimal minimum is considered “flat” if the test error changes less in its neighbourhood. Keskar et al. [11] and Chaudhari et al. [4] observe better generalization results for neural networks at flat minima. Gradient regularization has only been recently explored, that too from a numerical perspective for deep neural networks (DNNs) where it has been shown that fixed learning rates result in implicit gradient regularization  [22, 9, 1, 19].
@@
  • We propose a new line-search procedure called Constrained Gradient Descent (CGD) that performs gradient descent on a gradient regularized loss function. We also provide global linear convergence for CGD in Theorem 3.1.
@@ 2.1 Line-Search Methods
We focus on the situation where $\mathbf{x}$$^{*}$ is obtained using an iterative line-search procedure (for details refer [13, Ch. 3]). The steepest descent (gradient descent) method, Newton’s method, and quasi-Newton’s method are some examples of line-search methods. In an iterative algorithm, the key idea is to begin with an initial guess $\mathbf{x}_{0}$ for the minimizer and generate a sequence of vectors $\mathbf{x}_{0},\mathbf{x}_{1},\ldots$ until convergence (or until desired level of accuracy is achieved). In such algorithms, $\mathbf{x}_{k+1}$ is obtained using $\mathbf{x}_{k}$ using a pre-defined update rule.
@@ 2.2 Quasi-Newton Methods
We consider DFP and BFGS algorithms that update $G_{k}$ using symmetric rank-two updates at each descent step [3, 6]. The update equations for DFP and BFGS are given as follows (for more details refer [13, Ch. 6]):
$\textbf{(DFP) }G_{k+1}$ $=G_{k}+\frac{{\mathbf{s}}_{k}{\mathbf{s}}_{k}^{T}}{{\mathbf{y}}_{k}^{T}{\mathbf{s}}_{k}}-\frac{G_{k}{\mathbf{y}}_{k}{\mathbf{y}}_{k}^{T}G_{k}}{{\mathbf{y}}_{k}^{T}G_{k}{\mathbf{y}}_{k}}$
$\textbf{(BFGS) }G_{k+1}$ $=\left(I_{n}-\frac{{\mathbf{s}}_{k}{\mathbf{y}}_{k}^{T}}{{\mathbf{y}}_{k}^{T}{\mathbf{s}}_{k}}\right)G_{k}\left(I_{n}-\frac{{\mathbf{y}}_{k}{\mathbf{s}}_{k}^{T}}{{\mathbf{y}}_{k}^{T}{\mathbf{s}}_{k}}\right)+\frac{{\mathbf{s}}_{k}{\mathbf{s}}_{k}^{T}}{{\mathbf{y}}_{k}^{T}{\mathbf{s}}_{k}}$
@@
We can derive the corresponding Hessian approximation $\tilde{G}_{k}$ from the update steps for $G_{k}$ by applying the Sherman-Morrison-Woodbury formula [13, App. A]. These are given as follows:
$\textbf{(DFP) }\tilde{G}_{k+1}$ $=\left(I_{n}-\frac{{\mathbf{y}}_{k}{\mathbf{s}}_{k}^{T}}{{\mathbf{y}}_{k}^{T}{\mathbf{s}}_{k}}\right)\tilde{G}_{k}\left(I_{n}-\frac{{\mathbf{s}}_{k}{\mathbf{y}}_{k}^{T}}{{\mathbf{y}}_{k}^{T}{\mathbf{s}}_{k}}\right)+\frac{{\mathbf{y}}_{k}{\mathbf{y}}_{k}^{T}}{{\mathbf{y}}_{k}^{T}{\mathbf{s}}_{k}}$ (4)
$\textbf{(BFGS) }\tilde{G}_{k+1}$ $=\tilde{G}_{k}+\frac{{\mathbf{y}}_{k}{\mathbf{y}}_{k}^{T}}{{\mathbf{y}}_{k}^{T}{\mathbf{s}}_{k}}-\frac{\tilde{G}_{k}{\mathbf{s}}_{k}{\mathbf{s}}_{k}^{T}\tilde{G}_{k}}{{\mathbf{s}}_{k}^{T}\tilde{G}_{k}{\mathbf{s}}_{k}}$ (5)
@@ 3 Constrained Gradient Descent
Let $\mathcal{S}_{\hat{\mathbf{x}}}$ and $\mathcal{S}_{\mathbf{x}^{\star}}$ be the stationary points of $f({\mathbf{x}})$ and $g({\mathbf{x}})$ as defined in Equation 8. Then, $\mathcal{S}_{\hat{{\mathbf{x}}}}\subseteq\mathcal{S}_{{\mathbf{x}}^{*}}$. Furthermore, for any ${\mathbf{x}}^{*}\in\mathcal{S}_{{\mathbf{x}}^{*}}$ one of the following is true: (a) $^{\mathbf{x}}^{*}\in\mathcal{S}_{\hat{{\mathbf{x}}}}$ or (b) $\nabla f({\mathbf{x}}^{*})$ is an eigenvector of $H({\mathbf{x}}^{*})$ with the eigenvalue $\frac{1}{2\lambda}$.
@@
Two curves plotted over a shared x-axis: a blue curve labeled "Loss Landscape f(x)" and an orange curve labeled "Penalized Landscape g(x)". The penalized curve introduces more stationary points than the original. Black arrows mark the negative gradient directions on both curves, indicating the change in descent direction. Gradients are shown at the point x = -7.7, which is highlighted by a dashed vertical line.
@@
As a result, descent on the penalized function might actually cause ascent over the original loss function. For example in Fig. 2, observe that at x=7.7x=-7.7 (marked with the dashed line), steepest descent along $\nabla g({\mathbf{x}})$ direction would actually cause an ascent over the original loss function.
To fix this behaviour, we ensure that we only move along the direction $\nabla g({\mathbf{x}})$ when it is a descent direction. Therefore, we only move along $\nabla g({\mathbf{x}})$, if $\nabla f({\mathbf{x}})^{T}\nabla g({\mathbf{x}})>0$ (from Equation 2). Otherwise, we set $\lambda=0$ at this point and move along the $\nabla f({\mathbf{x}})$ direction. Note that this simple check also helps us to avoid stopping at artificially introduced stationary points. We summarize CGD in Algorithm 1.
@@
[…]
@@ 3.1 Global Linear Convergence of CGD
Substituting $\mathbf{x}_{k+1}\text{ and }\mathbf{x}_{k}$ in place of $\mathbf{y}\text{ and }\mathbf{x}$. And $\mathbf{x}_{k+1}-\mathbf{x}_{k}=-\alpha\nabla g_{k}$,
@@
3.1.1 CGD on Strongly Quadratic Function.
Next we investigate the rate of convergence of CGD on a strongly quadratic function as our objective. Such an analysis helps in identifying how penalization affects the condition number, which essentially controls the geometry of descent. For a matrix $A$, the condition number is defined as $\kappa(A)\triangleq\frac{\lambda_{\max}{(A)}}{\lambda_{\min}(A)}$. A higher condition number or ill conditioning implies a more zig-zagging behaviour and thus, slower convergence.
Let $f\left(\mathbf{x}\right)=\frac{1}{2}\mathbf{x}^{T}Q\mathbf{x}-\mathbf{b}^{T}\mathbf{x}$ where $Q\succ 0$. Let $l\triangleq\lambda_{\min}(Q)$ and $L\triangleq\lambda_{\max}(Q)$. From Equation 9 we know $\mathbf{x}_{k+1}=\mathbf{x}_{k}-\alpha B_{k}\nabla f_{k}$ where $B_{k}={I}+2\lambda Q$ and $\nabla f\left(\mathbf{x}_{k}\right)=Q\mathbf{x}_{k}-\mathbf{b}$. The error at each iterate is then given as:
$=\left\|{\left({I}-\alpha B_{k}Q\right)\left(\mathbf{x}_{k}-\mathbf{x}^{\ast}\right)+\alpha B_{k}\left(\mathbf{b}-Q\mathbf{x}^{\ast}\right)}\right\|\qquad\text{(on rearrangement) }$
$\therefore\left\|\mathbf{e}_{k+1}\right\|$ $\leq\left\|{{I}-\alpha B_{k}Q}\right\|\left\|\mathbf{e}_{k}\right\|$ (13)
Since I{I} and $Q$ commute, we have $(1+2\lambda l){I}\preccurlyeq B_{k}\preccurlyeq(1+2\lambda L){I}$. Also since $Q$ and $B_{k}$ are both symmetric positive-definite matrices and commute, we can say the following about $B_{k}Q$,
$\lambda_{\min}\left(B_{k}\right)\lambda_{\min}(Q)I\preccurlyeq B_{k}Q\preccurlyeq\lambda_{\max}\left(B_{k}\right)\lambda_{\max}(Q)I$
[…]
$\therefore\left\|{{I}-\alpha B_{k}Q}\right\|$ $\leq\max\left\{\left|1-\alpha\lambda_{\min}(B_{k})\lambda_{\min}(Q)\right|,\left|1-\alpha\lambda_{\max}(B_{k})\lambda_{\max}(Q)\right|\right\}$
$=\max\left\{\left|1-\alpha(1+2\lambda l)l\right|,\left|1-\alpha(1+2\lambda L)L\right|\right\}$
So in Equation 13,
$\left\|{\mathbf{x}_{k+1}-\mathbf{x}^{\ast}}\right\|$ $\leq\max\left\{\left|1-\alpha(1+2\lambda l)l\right|,\left|1-\alpha(1+2\lambda L)L\right|\right\}\left\|{\mathbf{x}_{k}-\mathbf{x}^{\ast}}\right\|$
$\leq\left(\max\left\{\left|1-\alpha(1+2\lambda l)l\right|,\left|1-\alpha(1+2\lambda L)L\right|\right\}\right)^{k}\left\|{\mathbf{x}_{0}-\mathbf{x}^{\ast}}\right\|$
Hence for CGD to converge,
$\max\left\{\left|1-\alpha(1+2\lambda l)l\right|,\left|1-\alpha(1+2\lambda L)L\right|\right\}<1,\ \forall k$
A suitable $\alpha$ that satisfies this condition is chosen to be
Both terms within the $\max(\cdot)$ achieve the same value at this $\alpha$.
$\therefore\left\|{\mathbf{x}_{k+1}-\mathbf{x}^{\ast}}\right\|$ $\leq\left(\frac{(1+2\lambda L)L-(1+2\lambda l)l}{(1+2\lambda L)L+(1+2\lambda l)l}\right)^{k}\left\|{\mathbf{x}_{0}-\mathbf{x}^{\ast}}\right\|$
where ${\kappa}=\kappa(B_{k})\kappa(Q)$.
Remark 1
Note that GD has a similar rate with the condition number there being $\kappa(Q)$ instead. And since $\kappa(B_{k})>0$, $\kappa>\kappa(Q)$. Thus, the condition number worsens while descending over the penalized landscape in comparison to the original objective. This indicates that our method would perform poorly for functions that have a high condition number and thus, take longer to converge.
@@ 3.3 CGD-QN: Quasi-Newton Variants of CGD
We consider quasi-Newton variants of CGD (CGD-QN) where a positive-definite approximation of the Hessian is maintained. The update step for CGD-QN is given as: ${\mathbf{x}}_{k+1}={\mathbf{x}}_{k}-\alpha\left(I_{n}+2\lambda_{k}\tilde{G}_{k}\right)\nabla f_{k}$ where $\tilde{G}_{k}$ is the Hessian approximation at step $k$, which is updated after every step. Particularly Equations 4 and 5 correspond to the updates of DFP and BFGS variants of CGD respectively. We summarize CGD-QN method in Algorithm 3.
@@
[…]
@@ 4 Numerical Experiments
We test CGD-FD and CGD-QN on synthetic test functions from Virtual Library of Simulation Experiments: Test Functions and Datasets  [^1] http://www.sfu.ca/˜ssurjano.
@@
Quadratic function 1010 0.4 0.01 18.89 97.91
Rotated hyper-ellipsoid function 55 0.5 0.01 15.94 82.76
Levy function 22 L(0.01, 0.1) 0.05 23.73 63.21
Branin function 22 0.07 0.01 37.53 87.07
Griewank function 22 40.0 0.01 0.01 0.08
Matyas function 22 10.0 0.01 1.83 34.40
Line plot showing drop in function value for CGD-FD vs GD on the Quadratic function. Both methods descend rapidly, with CGD-FD flattening earlier.
@@
Line plot showing drop in function value for CGD-FD vs GD on the Rotated hyper-ellipsoid function. Both methods descend rapidly, with CGD-FD flattening earlier.
@@
Line plot showing drop in function value for CGD-FD vs GD on the Levy function, with CGD-FD showing slightly faster early descent.
@@
Line plot showing drop in function value for CGD-FD vs GD on the Branin function. Both methods converge quickly, with CGD-FD slightly ahead.
@@
Line plot showing drop in function value for CGD-FD vs GD on the Griewank function. CGD-FD reaches a lower value and flattens earlier while GD decreases more gradually.
@@
Line plot showing drop in function value for CGD-FD vs GD on the Matyas function. CGD-FD reaches the minimum in fewer steps while GD decreases more gradually.
@@
Table 1 summarizes our findings and Fig. 3 visualizes function values vs gradient evaluations for the chosen test functions for experiments on CGD-FD. Observe that in the first step of optimizing Quadratic function (Fig. 3a), CGD-FD has a $97.91\%$ decrease in value compared to that of GD which only had an initial improvement of about $18.89\%$. Similar pattern can also be seen for other functions. This gives us the insight that CGD-FD penalizes the original function well enough to concentrate most of the decrease in value within the first few steps of the trajectory itself. Also, notice that in Griewank function (Fig. 3e) even though the initial step improvement is low ($0.08\%$), we still see a much rapid decrease in value within the first 10 steps compared to that of the GD trajectory. Thus, using CGD-FD with appropriate penalization can provide for great boosts in function value well within the starting iterates of the optimization procedure.
Line plot showing drop in function value for CGD-QN vs QN methods on the Zakharov function. CGD-QN methods converge faster.
@@
Line plot showing drop in function value for CGD-QN vs QN methods on the Drop-Wave function. Both CGD-QN variants reach lower values earlier, while their vanilla counterparts plateau.
@@
Line plot showing drop in function value for CGD-QN vs QN methods on the EggHolder function. Both CGD-QN methods show substantial improvement while their vanilla counterparts remain flat.
@@
4.0.1 Experiments for CGD-QN.
We ran CGD-BFGS and CGD-DFP methods against BFGS and DFP for $T=40$. The function values vs iterations are visualized in Fig. 4. We observe that for most suitable choices of $\lambda$, CGD-DFP and CGD-BFGS exhibit similar behaviour and hence, their trajectories coincide. Another thing to observe is that in CGD-QN methods, most improvements aren’t concentrated in the initial steps since Hessian approximation is still not very great here. Instead the improvements are more gradual and appear over further steps. For example, in Zakharov function (Fig. 4a), we can observe that all methods follow the same decrease in function value for the first 1–2 steps and then the CGD-QN methods start to drop more in value. In EggHolder function (Fig. 4c) this effect is most pronounced where the improvements are very gradual but providing with better results than their vanilla counterparts.